### v1

# VisualLens: Personalization through Visual History

Abstract
We hypothesize that a user’s visual history with images reflecting their daily life, offers valuable insights into their interests and preferences, and can be leveraged for personalization. Among the many challenges to achieve this goal, the foremost is the diversity and noises in the visual history, containing images not necessarily related to a recommendation task, not necessarily reflecting the user’s interest, or even not necessarily preference-relevant. Existing recommendation systems either rely on task-specific user interaction logs, such as online shopping history for shopping recommendations, or focus on text signals. We propose a novel approach, VisualLens, that extracts, filters, and refines image representations, and leverages these signals for personalization. We created two new benchmarks with task-agnostic visual histories, and show that our method improves over state-of-the-art recommendations by 5-10% on Hit@3, and improves over GPT-4o by 2-5%. Our approach paves the way for personalized recommendations in scenarios where traditional methods fail.
[…]
Wang Bill Zhu at
1 Introduction
Figure 1: VisualLens leverages a user’s task-agnostic visual history to provide personalized recommendations. Our method outperforms GPT-4o by 1.6$\sim$4.6on Hit@3.
Imagine a personal assistant observing what you do in your daily life. With her keen insight, she can make informed guesses about what you may enjoy or find intriguing. When you ask for recommendations on anything from restaurants and activities to movies, books, and products, based on her in-depth understanding of you she will provide suggestions tailored specifically to your tastes.
Although the idea seems straightforward, a comprehensive personal assistant that can make recommendations across various aspects of life has yet to become a reality. Most existing recommendation systems are limited to specific domains and rely on interaction history within the particular verticals (Tan and Jiang, 2023; Papadakis et al., 2023). For instance, an e-commerce platform may suggest products based on your past purchases, but it will not take into account your dining habits or interests outside of shopping. Similarly, a restaurant booking system may recommend eateries based on our previous reservations, but it will not know your new interest in cooking or your recent travels. A video streaming service may propose movies based on your viewing history, but it will not be aware of your cultural curiosities.
Recent advances in personalized conversation systems have utilized task-agnostic conversation logs to provide personalized answers (Harper and Konstan, 19; Ni et al., 2019; Liu et al., 2023). However, these approaches predominantly rely on textual data.
We hypothesize that a user’s visual history, comprising images reflecting their daily life, offers valuable insights into their interests and preferences. This work aims to investigate whether such a visual history can enable a recommendation system to better understand a user’s individual preferences, and deliver more personalized recommendations.
There are many challenges in achieving this goal. First, although we envision a future system, akin to Vannevar Bush’s Memex (Bush, 1945), that can record a user’s visual history with her consent, such a system remains elusive because of hardware limitations such as battery life, thermal constraints, and storage capacity. Second, a user’s visual history is inherently diverse and noisy, containing images not directly related to a specific recommendation ask, entities not accurately reflecting the user’s preferences (e.g., visiting a museum without enjoying it), and elements not preference-indicative (e.g., background objects like trash cans). Lastly, since many recommendation tasks require instant responses, we need to develop a low-latency solution.
In this paper, we propose a novel approach VisualLens, as a first step towards harnessing a user’s visual history for personalized recommendation. Our first contribution tackles the challenge of history availability by leveraging photos taken by a user. Unlike extensive visual logs, the photos a user takes require significantly less storage, oftentimes are readily available in reviews and social media posts, and convey more meaningful signals about a user’s interests and preferences. To facilitate evaluation, we created two benchmarks—Google Review-V and Yelp-V—using publicly available photos, providing a foundation for personalization assessments.
Our second contribution is a recommendation solution that leverages a user’s visual history to improve personalization. At the core of our solution is a set of models that effectively extract the essential signals from the visual history to infer the user’s preferences. First, given a recommendation request, we selectively retrieve only the most relevant images, thereby reducing unnecessary noises and distractions. Second, to effectively capture the signals a photo may convey, we use not only visual embeddings but also text captions and aspect words extracted from the image. Third, we adopt an iterative refinement process that continually improves aspect extraction to better reflect user interests and inform recommendations. By jointly training aspect word extraction and prediction tasks within a unified model, VisualLens not only reduces the overall parameter size, but also enhances the model’s capability to understand and utilize the visual history for accurate recommendation.
Our third contribution is a pipeline that applies multiple optimizations to accelerate recommendation generation. To minimize online processing time, we pre-generate image captions and extract aspects offline. Furthermore, we introduce a grid-based approach that aggregates multiple images into a $d\times d$ grid, enabling simultaneous processing of multiple images and significantly reducing computation overhead.
[…]
2 Related Works
Recommendation system with large language models.
Large Language Models (LLMs) have demonstrated strong potential in recommendation systems with their advanced language processing capabilities (Tan and Jiang, 2023). On item-based recommendation, studies such as LLM4RS (Dai et al., 2023), LLMRank (Hou et al., 2024), CLLM4Rec (Zhu et al., 2024b), P5 (Geng et al., 2022), and Sanner et al. (2023) explored various LLM prompt and bootstrapping strategies, showing competitive performance, especially in cold-start scenarios. Generative recommendation with open-domain prompts is explored by GenRec (Ji et al., 2023), though fine-tuning remains crucial. Fine-tuning approaches include personalized aspect extraction (Li et al., 2023c), and multitasking on LLaMa models (Yang et al., 2023). Retrieval-enhanced models, such as ReLLa, improve recommendation by retrieving relevant behavior sequences (Lin et al., 2024). Instruction-tuning and graph augmentation approaches are explored in InstructRec (Zhang et al., 2023), LLMRec (Wei et al., 2024b), and LKPNR (Hao et al., 2023). LLMs also work well with content-based recommendation, such as Li et al. (2023a) on prompting, and ONCE (Liu et al., 2023), LoID (Xu et al., 2024) on fine-tuning.
Multimodal large language models.
Multimodal LLMs are becoming increasingly powerful in processing images and videos and in generating human-like text, exemplified by models like GPT-4 family (OpenAI, 2024), Claude 3.5 family (Anthropic, 2024), Gemini and PaliGemma (Google, 2023; Beyer et al., 2024), LLaVA model family (Liu et al., 2024), and Llama 3 Vision models (Meta, 2024). However, they still suffer from a strong language prior (Fu et al., 2024a; Parcalabescu and Frank, 2024), generalization (Zhu et al., 2022) and hallucinations (Fu et al., 2024b), or require heavy text transcription (Zhu et al., 2024a). The extent to which they can understand user history, in analogy to LLM recommendation systems, is still unclear.
Multimodal recommendation systems.
Multimodal recommendation systems leverage multiple data types, such as text and images, to improve recommendation relevance and personalization. Before the LLM era, Lee and Abu-El-Haija (2017) proposed systems for content-only video recommendation using similarity learning. PC2(Yan et al., 2023) develop an LLM model that provides multimodal explanations for recommendations. On modalities beyond image and text, LSVCR (Zheng et al., 2024) built a joint recommendation system integrating comments with video items (Davidson et al., 2010; Du et al., 2020; Gao et al., 2017; Zhou et al., 2024). The current state-of-the-art image-text recommendation, UniMP (Wei et al., 2024a), extended single-task multimodal personalization (He and McAuley, 2016; Tang et al., 2020; Wei et al., 2019, 2023) on multitask website-based shopping. Methods such as RecFormer (Li et al., 2023b) and UniSRec (Hou et al., 2022) convert images to short captions to utilize text-only models. However, none of these methods consider the scenario of personalized recommendation from the user’s visual history. In this work, we present a novel method VisualLens leveraging multimodal LLM to achieve personalization with visual history.
Figure 2: VisualLens inference pipeline: the offline process augments images in the visual history with captions and aspect words; the runtime recommendation process retrieves relevant images, generate user profile accordingly, and then predict candidate preferences.
3 Problem Definition
Consider a recommendation QA (question answering) task, where the user asks a recommendation question $q$, and the recommender answers $q$ with a ranked list of items. Good recommenders shall rank the items that the user is more likely to be interested in or willing to try early in the list.
The recommender is facilitated with a task-agnostic visual history ${\cal H}_{u}$ for each user $u$, which contains a series of photos, taken or posted by the user, not necessarily related to $q$. We state three assumptions to allow generalization. First, the photos may not be directly relevant to the question. Second, an image may not be associated with any candidate item, and even so, the candidate ID is not given. Third, a photo does not necessarily imply strong preferences. Figure 1 shows an example question, visual history, and candidates.
To simplify the problem, we assume a candidate retriever exists to retrieve all candidates that satisfy the user’s question. Each candidate $s$ is represented with a $(x_{s},{\cal I}_{s})$ pair, where $x_{s}$ is the name and text descriptions, and ${\mathcal{I}}_{s}$ is the image set for the item.
Traditional recommendation setting considers two types of signals. The first is a task-specific set of items in the candidate set, which captures user interest or at least user history. The second is a set of user-specific attributes such as the user’s age, gender, and interests. This paper focuses on task-agnostic visual history and leaves integration of these traditional signals for future extensions.
4 Recommendation Method
We start with describing the VisualLens solutions. Our solution contains two parts: offline history augmentation, and runtime recommendation generation.
  • Offline history augmentation4.1) augments each image in the visual history with captions and aspect words extracted from the image. Figure 2 shows example captions and aspect words of the images in the visual history.
  • Runtime recommendation generation4.2) answers a recommendation question $q$ in three steps. First, the history retrieval step retrieves only images relevant to $q$, since the visual history can be diverse and not all photos are relevant to every query. Second, the preference profiling step uses the retrieved images and their augmented captions and aspects to generate the user’s preference profile. Third, the candidate matching step matches the user’s preference profile with each candidate, to generate the confidence score for each candidate for ranking. VisualLens utilizes the same backbone model with parameter $\theta$ for aspect generation and candidate matching.
4.1 Offline history augmentation
To best capture various signals in an image, we encode it to visual embeddings and augment it with a caption and a set of aspect words. As we show in the experimental ablation study (Table 6.3), these augmentations allow us to leverage the strengths of various encoding techniques and considerably improve the recommendation quality. These augmentations can be used directly at recommendation time to reduce runtime latency.
Image encoding.
We encode each image in the visual history using the CLIP ViT-L/14@336px model (Radford et al., 2021). This embedding will be directly leveraged in image retrieval in the history retrieval step at runtime.
Caption generation.
We generate image captions using a frozen LLaVA-v1.6 8B model (Liu et al., 2024). We prompted the model to generate captions within 30 words, such that the captions are succinct and short of hallucinations.
Aspect word generation.
Aspect words are descriptive terms that capture essential features or attributes of an image (e.g., dome, balcony, plants). Aspect words provide crucial details about the image. We prompted the backbone model to list aspect words for an image. We do not constrain the number of extracted aspect words since each image may have varying numbers of aspect words essential for final preference prediction. We will describe in Section 5.3 how we finetune the model to generate aspect words that are most useful for recommendation tasks.
4.2 Runtime recommendation generation
History retrieval.
Given a query $q$ and a user’s visual history ${\cal H}_{u}$, VisualLens first retrieves images that are related to $q$, denoted by $\mathcal{I}_{u,q}$. We choose up to $w$ images to cap the number of images we process at runtime, and ensure that only the most contextually relevant images are retained for further processing, thereby reducing noise.
In general, we can use any image retrieval method such as DELG (Cao et al., 2020). Here, we present a method for categorical recommendations such as restaurants and museums, popular in recommendation tasks. For each category $c$, we randomly select a set of candidate items in the category and average the visual embeddings of their images as the image embedding of the category, denoted by $\mathbf{v}_{c}$. Specifically, the category embedding is calculated as:
$\mathbf{v}_{c}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{c}^{(j)}$
where $n$ is the number of candidates, and $\mathbf{v}_{c}^{(j)}$ indicates the visual embedding of the $j$-th item image in category $c$. The retrieval step measures the cosine similarity between the visual embedding $\mathbf{v}_{i}$ of each image $i\in{\cal H}$ in the user’s history and the image embedding $\mathbf{v}_{c}$ of the relevant category $c$. We then select the top-$w$ images based on the cosine similarity scores.
Preference profiling
. Given a set of retrieved images $\mathcal{I}_{u,q}$, VisualLens then generates the user’s preference profile relevant to the images, their captions, and aspect words.
A critical part of this step is image encoding. Even after retrieval, the number of images $w$ is still large. Most multimodal LLMs allow context windows of limited sizes, constraining the number of images we can process. For example, for an input image of resolution $896\times 896$, the PaliGemma model would generate an embedding of up to 4,096 tokens. A typical LLM with a context window of 8,192 tokens can take at most 2 images.
We propose to group relevant images $\mathcal{I}_{u,q}$ into a $d\times d$ grid, where $d^{2}=w$, and treat all images in the grid as a single image. If we have retrieved fewer than $w$ images, we pad with a black background. Let $h$ be the maximum available resolution in a multimodal LLM. The gridify process $G$ takes the $d\times d$ grid and generates an image of fixed size $\mathbb{R}^{h\times h\times 3}$. Additionally, we number each image from 1 to $d^{2}$ to ensure the images are grounded to the corresponding caption and aspect words in the input to the candidate matching.
We denote a user $u$’s preference profile on question $q$ by $(i_{u,q},x_{u,q})$, where $i_{u,q}$ denotes the gridified image, and $x_{u,q}$ denotes the concatenated captions and aspect words of relevant images.
Candidate matching.
Finally, VisualLens takes the user’s preference profile $(i_{u,q},x_{u,q})$ and a set of candidates, each represented by $(x_{s},{\mathcal{I}}_{s})$, and generates the matching score for each candidate, which will then be used for ranking. This is achieved by prompting the multimodal candidate predictor, where we packed the user profile and candidates to the prompt through the image channel and the text channel separately (see the template in Appendix C).
5 Iterative Refinement and Joint Training
VisualLens requires LLMs for image encoding, image captioning, aspect word generation, and final preference-candidate matching. Out-of-box LLMs do not perform well, and we conduct fine-tuning in three steps.
  • First, to better encode the $d\times d$ grid of images, we fine-tune image encoding with multi-image caption tasks.
  • Second, since aspect generation is not a standard multimodal task, we iteratively refine the aspect word extractor, ensuring that the keywords precisely capture the aspects of the image in a suitable granularity.
  • Third, to best leverage personal profile signals including image embeddings, captions, and aspect words, we finetune the candidate matching model. This is done through joint multitask training with the aspect words generation.
5.1 Multi-image caption pretraining
To facilitate the model to ground to each grid faithfully, we perform a LoRA continual pretraining on dense captions. We adopt the dense captioning dataset DOCCI dataset (Onoe et al., 2024), which contains over 15,000 images and their corresponding dense captions. Each time, we randomly sample $w$ images =$\mathcal{I}=\{i_{1},\cdots,i_{w}\}$ and their corresponding caption $\mathcal{C}=\{x_{1},\cdots,x_{w}\}$, and then we construct a gridified input image $G(\mathcal{I})$ and a target output text description $T(\mathcal{C})=$Image 1: $x_{1}$, $\cdots$, Image w: $x_{w}$”. Then we LoRA finetune the pretrained backbone model (e.g., MiniCPM-V2.5) on all image-caption pairs $\{G(\mathcal{I}),T(\mathcal{C})\}$ so that the model is able to process the gridified user history grid by grid. We then use the continual pretrained model as the starting point to apply joint training described in §5.3.
5.2 Iterative aspect word refinement
Unlike image captioning, aspect word generation is not a standard multimodal task and lacks the extensive pretraining data typically available. Hence, with zero-shot prompting, the generated aspect words have a large quality gap across different images, and the extracted aspects may not indicate user preferences.
To finetune aspect word generation, we first generate the training data. For each image $i$, we start from an initial set of aspect words, denoted as $\mathcal{W}_{i}^{(0)}$, which is generated by LLaVA-v1.6. In the $j^{\text{th}}$ round, we prompt a separate Llama-3.1 70B model with $\mathcal{W}_{i}^{(j-1)}$, candidates, and ground truths, and ask it to select useful aspect words that are helpful in ground truth prediction, which constitute $\mathcal{W}_{i}^{(j)}$. This refinement process continues for several rounds, and the iterations allow for converging extracted aspect words toward a more accurate and relevant subset. Empirically, we observe that the refinement converges after approximately 4 rounds, and denote ${}^{\textnormal{th}}$ refined aspect word set $\mathcal{W}_{i}^{(4)}$ as $\mathcal{W}_{i}$, which serves as the training target.
The backbone model with parameter $\theta$ is finetuned to optimize the cross-entropy (CE) loss over all images $\mathcal{I}$,
$\mathcal{L}_{\text{asp}}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\textsc{CE}(\mathcal{W}_{i},p_{\theta}(x_{\text{asp}},i)),$ (1)
where $x_{\text{asp}}$ is the prompt for aspect words generation.
5.3 Joint training of aspect word generation and candidate matching
To take advantage of multitask training in multimodal recommendation (Wei et al., 2024a), we jointly train the aspect word generator and the candidate predictor on the backbone model. This joint training strategy allows the model to simultaneously learn to identify useful aspect words and make accurate predictions, thus improving overall performance.
The joint loss function balances aspect word generation and candidate matching with a weighting factor $\lambda$, where the candidate matching is optimized with binary cross-entropy (BCE) loss to handle multiple ground truth labels.
$\mathcal{L}_{\text{pred}}=\frac{1}{N}\sum_{j=1}^{N}\textsc{BCE}(\mathcal{S}_{j},p_{\theta}(x_{\text{pred},j},i_{u_{j},q_{j}})),$ (2)
$\mathcal{L}_{\text{joint}}=\mathcal{L}_{\text{asp}}+\lambda\mathcal{L}_{\text{pred}},$ (3)
where $u_{j},q_{j},\mathcal{S}_{j}$ is the user, the question, and the ground truth set of candidates of the $j$-th example. The text prompt $x_{\text{pred},j}$ consists of the question $q_{j}$, the text user preference profiles $x_{u_{j},q_{j}}$, and candidates. We LoRA finetune the model under the joint loss $\mathcal{L}_{\text{joint}}$.
{NiceTabular}
[…]
GR-V 15.69M 2K 200K 66 157.0 2.7 43.1
Yelp-V 4.12M 2K 100K 35 263.6 8.2 66.7
Table 1: Dataset statistics of Google Review-Vision (GR-V) and Yelp-Vision (Yelp-V).
{NiceTabular}
[…]
Modality Size Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10 MRR
 Naive baselines
Random - - 7.6 21.0 55.0 21.2 13.0 33.6 72.7 30.0
Rank by rating - - 3.9 15.8 55.5 17.7 8.7 28.0 72.3 25.9
 Fine-tuned models
UniMP (Wei et al., 2024a) T + I 3B 13.8 34.1 73.0 30.5 22.4 48.5 85.0 38.3
Llama-3.1-8B-Instruct (Meta, 2024) T 8B 15.8 36.3 77.2 32.9 24.1 52.2 88.5 39.6
MiniCPM-V2.5 (Yao et al., 2024b) T + I 8B 16.1 36.4 78.4 33.2 24.8 53.0 89.3 40.3
 Direct inference
Llama-3.1-70B-Instruct (Meta, 2024) T 70B 16.2 35.9 75.7 33.1 25.2 53.2 88.5 40.6
GPT-4o (OpenAI, 2024) T + I - 17.1 37.3 80.1 34.3 26.1 54.5 90.5 41.7
 Our method
VisualLens
(PaliGemma) T + I 3B 16.7 36.3 77.1 33.5 27.8 58.8 90.4 44.3
VisualLens (MiniCPM-V2.5) T + I 8B 18.5 38.9 82.3 35.4 28.3 59.1 91.0 44.9
Human annotations - - 22.0 45.0 - - 36.0 66.0 - -
Table 2: Hit rates and MRRs of VisualLens vs. multiple baselines on Google Review-V and Yelp-V. The result shows (a) VisualLens outperforms other baselines, though has a gap with the human oracle; (b) model size greatly affects the performance; (c) simply rank by rating is a worse design than the random baseline.
6 Benchmarks and Experiments Setups
6.1 Benchmark creation
To the best of our knowledge, there is no existing benchmark (Harper and Konstan, 19; Ni et al., 2019; Wu et al., 2020; Wan and McAuley, 2018; Salemi et al., 2023) to evaluate personalization with task-agnostic visual history. We created two benchmarks, Google Review-V and Yelp-V, leveraging publicly available data from Google Local Review (Li et al., 2022) and Yelp (Asghar, 2016).
User logs:
For each user in the two datasets, we take the list of reviews in chronological order. Each review is associated with a business name, categories and description. In Google Review-V, each review is associated with a few photos, used for image logs. Yelp-V does not associate a review with photos, so we randomly subsample one-third of the store profile pictures, such that different reviews for the same business can be associated with different images.
Questions and visual history:
We consider a special type of questions, category recommendations, like "Recommend a nearby museum". Such questions are both popular in real applications, and hard as there are many candidates satisfying the constraint. We remove small categories and most ambiguous categories, such as “place”, and “spot”.
For each review $r$ regarding a business of category $c$, we create a question asking to recommend businesses in category $c$. We take all (and only) photos in the reviews before $r$ to generate the visual history. We consider the visual history task-agnostic since the categories are highly diverse (see Figure 4, and the photos are quite diverse too (e.g., a park photo to illustrate happiness mentioned in the review). We filter instances where the history is too short ($<$$10$) or does not contain the questioned category.
Candidates and ground truths:
For a review $r$ we take all reviews starting from $r$ to generate candidates and ground truths. To be realistic, we consider only nearby businesses of the same category as the candidate set, and the number of candidates is a random number in $[30,100]$. Candidates that also appear in the user’s future reviews are considered as ground truths. To avoid falling into a classification problem, we filter examples with only 1 ground truth in Google Review-V and fewer than 5 in Yelp-V.
Summary:
Table 5.3 gives the benchmark statistics. The ratio of an average number of candidates and that of ground truths is 16:1 for Google Review-V and 8:1 for Yelp-V, showing the difficulties of recommendation. By default, the train, dev, and test data have disjoint users. We discuss other splitting in Table 6.3.
6.2 Evaluation measures
We use two metrics to evaluate recommendation quality.
Hit@$\mathbf{k}$.
[…]
Mean Reciprocal Rank (MRR).
$\text{MRR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{\text{rank}(r_{i})}$ measures the ranking quality by averaging the reciprocal ranks of the first relevant item for each example. The MRR ranges from $1/S$ to $1$, where $S$ is the number of candidates.
6.3 Implementation and baselines
We ran VisualLens with two backbone models, a smaller 3B model PaliGemma and a larger 8B model MiniCPM-V2.5. We used an image grid of $8\times 8$, and a loss weighting factor $\lambda=2$ in Eqn. 3. We compared VisualLens with multiple other solutions.
  • Baselines: Randomly rank or select top-$k$ ratings.
  • Fine-tuned models: We fine-tuned three state-of-the-art solutions: multimodal personalization model UniMP (Wei et al., 2024a) (RedPajama 3B model), with the adaptation to replace their item image and attributes with image tokens in the visual history; Llama-3.1-8B-Instruct (Meta, 2024) with text-only user preference profiles; and MiniCPM-V2.5 8B (Yao et al., 2024a) with image profiles.
  • Direct inference: We compared with two out-of-box models: Llama-3.1-70B-Instruct (Meta, 2024) with text-only preference profiles; GPT-4o (OpenAI, 2024) with multimodal profiles. To control the API cost, we subsample 1k instances from the test sets.
  • Human annotation: Finally, we subsampled 50 examples from the test set for human annotation.
{NiceTabular}
[…]
Asp. Cap. Img. Iter. Joint Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10 MRR
1 ✓ ✓ ✓ ✓ ✓ ✓ 16.7 36.3 77.1 33.5 27.8 58.8 90.4 44.3
2 ✓ ✓ ✓ ✓ ✓ 16.1 35.8 76.2 33.0 27.2 57.9 88.9 43.3
3 ✓ ✓ ✓ ✓ 15.7 35.2 75.4 32.5 26.9 57.5 88.2 42.9
4 ✓ ✓ ✓ 15.2 34.7 74.2 31.9 25.8 55.3 86.1 41.2
5 ✓ ✓ ✓ 14.8 33.9 73.0 31.2 25.0 53.9 84.9 40.4
6 ✓ ✓ 13.5 32.5 71.9 29.6 22.0 48.2 83.6 38.8
7 ✓ ✓ ✓ 11.5 27.9 67.3 25.9 20.1 45.7 81.7 36.8
Table 3: Ablation study on PaliGemma. Different components of VisualLens model: joint training (Joint), iterative refinement (Iter), aspect words (Asp.), captions (Cap.), image embedding (Img.), and relevant image retrieval (Ret.). MRR difference over 0.4 has a p-value$<$0.04.
Figure 3: (a) MRR distribution over number of candidates, (b) MRR distribution over number of images. Both are on the User ID test set. We find (1) MRR converges when number of candidates exceeds 50; (2) MRR increases and flattens after reaching $\sim$100 images.
{NiceTabular}
[…]
LongHis Category User ID LongHis Category User ID
VisualLens (PaliGemma) 35.9 34.9 33.5 46.6 45.2 44.3
VisualLens (MiniCPM-V2.5) 38.0 37.1 35.4 47.2 45.5 44.9
Table 4: Transferability: MRR of VisualLens models to different test setups. LongHis: train until a certain timestamp and test afterwards. Category: held-out a set of categories for testing per user. Use ID: train and test set share no common user ID.
Figure 4: (a) MRR distribution over categories on Google Review-V, (b) MRR distribution over categories on Yelp-V. We find (1) the performance per category is loosely correlated with number of training data; (2) when a category is more general and less ambiguous, the performance on the category is better.
7 Results and Analysis
We conducted experiments to answer four questions:
  • Q1: Can we effectively leverage a user’s visual history to improve personalization?
  • Q2: How does each element of VisualLens contribute to the recommendation quality?
  • Q3: Can VisualLens transfer across users, unseen categories and longer history?
  • Q4: What is the robustness of VisualLens?
7.1 Recommendation effectiveness (Q1)
VisualLens significantly outperforms baselines.
Our first observation from Table 5.3 is that all recommendation models that leverage the visual history significantly outperform baseline solutions on all metrics. In particular, the best version of VisualLens improves Hit@3 over Random by 18% on Google Review-V and by 26% on Yelp-V, and even more over Rank by rating (apparently, overall ratings do not cater to specific users). There is still a gap between VisualLens and human annotations, but comparing w. Random, it fills $\sim$75of the gaps for hit@3 on both datasets.
VisualLens outperforms state-of-the-art solutions.
VisualLens outperforms UniMP with the same number of trainable parameters. With the 8B MiniCPM-V2.5 backbone, VisualLens outperforms MiniCPM-V2.5 itself by 2.5% on Hit@3 on Google Review-V, and by 6% on Yelp-V, and we observe a similar trend for the 3B models. Even comparing with significantly larger models, including Llama-3.1-70B-Instruct and GPT-4o without fine-tuning, VisualLens 7B improves by 1.6-5.6% on Hit@3.
7.2 Ablation studies (Q2)
We evaluate the usefulness of each component in VisualLens in Table 6.3, with PaliGemma as the backbone.
  1. 1.
    History retrieval significantly improves the results, and can reduce Hit@3 by 7% on Google Review-V and by 12% on Yelp-V (#7 vs. #3).
  2. 2.
    All three representations of the images (embedding, caption, aspects) play an important role. Removing captions and aspect words can reduce Hit@3 by 3% on Google Review-V and by 9% on Yelp-V, even without fine-tuning (#6 vs. #3). Between the two, aspect words play a more important role than captions (#5 vs. #4).
  3. 3.
    Both iterative training and joint multi-task training improve the recommendation quality. Removing both of them lowers Hit@3 by 1%+ on both data sets (#3 vs #1).
7.3 Transferability (Q3)
We tested transferability over users (default setting), over categories, and over different (longer) history lengths. Table 6.3 compares the MRR of VisualLens on both benchmarks, showing good transferability. MRR is highest when applied to longer history, with the effectiveness of history retrieval. Transferability is higher across categories than across users, both demonstrating the promise of leveraging task-agnostic signals, and illustrating slightly more challenges to transfer between users of different interest patterns.
7.4 Robustness (Q4)
Finally, we conducted a few robustness tests.
  1. 1.
    Figure 3(a) shows that as the number of candidates grows, the recommendation becomes harder and MRR gets lower. However, when the number of candidates exceeds 50, MRR converges.
  2. 2.
    Figure 3(b) shows that as the history grows, MRR increases and flattens after reaching $\sim$100 images. This trend is related to our grid size $8\times 8$, since smaller than 64 images would not leverage all spaces in the grid. On the other hand, flat MRR after 100 images shows robustness against history noises with the retrieval step.
  3. 3.
    Finally, Figure 4 plots MRR for different categories. There are a few factors that can affect recommendation quality. First and foremost, ambiguous categories like “area”, “station”, and “market” get the lowest MRR in Google Review-V. Second, general categories (e.g., “museum”, “hotel”) with bigger category sizes obtain higher MRR than specific ones (e.g., “historical landmark”). Third, transfer learning happens between neighbor categories; for example, “deli” and “takeout” achieve the top-$2$ and top-$3$ performance with less training data, since they are similar to the largest category “restaurant”.
8 Conclusion and Discussion
In this paper, we proposed a novel approach VisualLens for personalized recommendation using task-agnostic visual history. We advanced multimodal LLMs to extract meaningful signals from images to serve as proxies of user preferences. We created two benchmarks, Google Review-V and Yelp-V, to evaluate VisualLens, affirming the efficacy and robustness of VisualLens.
VisualLens lends itself as a first step towards recommendation systems leveraging task-agnostic visual history. There are still many future directions to explore. First, we can integrate the VisualLens solution with other accessible information, such as the date and location each image is taken, the specific fine-granularity entities (e.g., particular products) recognized from the images, and other user profiles. Second, we can extend the recommendation problems studied in this paper with broader QA tasks. Finally, we shall study how to guarantee privacy in the training with techniques like federated learning.
Acknowledgements
The robot icon used in Figure 1 is from https://www.flaticon.com/free-icon/robot_2432846. The padlocks icons used in Figure 2 are from https://www.flaticon.com/free-icon/lock_996365 and https://www.flaticon.com/free-icon/padlock_535143. The illustration images in Figures 1, 2 are from DOCCI (Onoe et al., 2024).
A Benchmark Creation Details
A.1 Algorithm for candidate generation
We list the complete candidate and ground truth sets generation algorithm in Algorithm 1, where the function nearest($\mathcal{G},\text{loc},m$) returns the $m$ nearest businesses of around certain location loc based on the graph $\mathcal{G}$. The function unique_name($\mathcal{S}$) removes the businesses in the set $\mathcal{S}$ with redundant names.
In both Google Review-V and Yelp-V, ${rc}_{\text{min}}=30$, ${rc}_{\text{max}}=100$, ${fc}_{\text{min}}=10$.
Algorithm 1 Candidate and Ground Truth Sets Generation
Input: Business geographics graph $\mathcal{G}=(\cal{V},\cal{E})$, Visit set $\cal{B}$ of a user, Category $c$, Minimum random candidate count ${rc}_{\text{min}}$, Maximum random candidate count ${rc}_{\text{max}}$, Minimum final candidate count ${fc}_{\text{min}}$, Minimum ground truth count ${gtc}_{\text{min}}$.
Output: Candidate sets $\mathcal{S}^{\text{CD}}_{1..n}$, Ground truth sets $\mathcal{S}^{\text{GT}}_{1..n}$.
# Initialization
for each business $b\in\cal{B}$ do
     flag $f_{b}\leftarrow unselected$
end for
$count\leftarrow 0$
# Main algorithm
for each business $b\in\cal{B}$ do
     if $b$ is in category $c$ and $f_{b}=unselected$ then
         $count\leftarrow count+1$
         Candidate count $m\leftarrow\text{random}(rc_{\text{min}},rc_{\text{max}})$
         Candidate set $\mathcal{S}^{\text{CD}}_{count}\leftarrow\text{nearest}(\mathcal{G},b_{\text{loc}},m)$
         Candidate set $\mathcal{S}^{\text{CD}}_{count}\leftarrow\text{unique\_name}(\mathcal{S}^{\text{CD}}_{count})$
         Ground truth set $\mathcal{S}^{\text{GT}}_{count}\leftarrow\mathcal{S}^{\text{CD}}_{count}\cap\cal{B}$
         if $|\mathcal{S}^{\text{CD}}_{count}|<fc_{\text{min}}$ or $|\mathcal{S}^{\text{GT}}_{count}|<gtc_{\text{min}}$ then
              $count\leftarrow count-1$
              continue
         end if
         for each business $b^{\prime}\in\mathcal{S}^{\text{GT}}_{count}$ do
              flag $f_{b^{\prime}}\leftarrow selected$
         end for
     end if
end for
return $(\mathcal{S}^{\text{CD}}_{1..count},\mathcal{S}^{\text{GT}}_{1..count})$
Figure 5: The Google Review-Vision (Google Review-V) training data consists of 66 categories.
A.2 Processing and filtering
Category removal and merge.
For Google Review-V and Yelp-V, the categories are selected as the last word of the annotated tags in the business. However, some category requires multiple words to express its meaning, such as “tourist attraction”, “steak house”, “historical landmark", “nature preserve”, etc. We select and keep these multi-word categories.
In Google Review-V, we remove smaller categories with less than 10k occurrence in the dataset. Then, we remove ambiguous or non-differentiable categories, including “shop”, “store”, “complex”, “service”, “company”, “supplier”, “caterer”, “agency”, “center”, “organization”, “attraction”, “house”, “mall”, “landmark”, “wash”, “course”, “preserve”, “alley”, “groomer”, “field”, “peak”, “venue”, “delivery”, “dealer”, “lounge”, “office”, “arcade”, “court”, “spot”, “stop”, “maintenance”, “trainer”, “wholesaler”, “planner”, “place”, “facility”, “school”, “stand”, “range”, “consultant”, “designer”, “veterinarian”, “ground”, “contractor”, “manufacturer”, “studio”, “point”, “lot”.
In Yelp-V, as the category is very centralized, we remove smaller categories with less than 50k occurrence in the dataset. Then, we remove ambiguous or non-differentiable categories, including“planning”, “nightlife”, “services”, “wings”, “arts”, “dogs”, “tacos”, “caribbean”, “beer”, “spirits”, “wine”, “venues”, “fusion”, “entertainment”, “southern”, “spaces”, “lounges”, “breweries”, “shopping”, “smoothies”, “flavor”, “plates”, “eastern”, “tex-mex”, “shop”, “noodles”, “markets”, “market”, “donuts”, “gelato”, “sum”, “veggies”, “fruits”, “trucks”, “bagels”, “cheesesteaks”, “clubs”, “cuban”, “ramen”, “life”, “roasteries”, “stands”, “brewpubs”, “gluten-free”, “gardens”, “travel”.
A.3 Training data distribution
We list the training data distribution of Google Review-V in Figure 5.
B More Related Works
B.1 Traditional recommendation methods.
Before the LLM era, there are also a line of work for recommendation with other networks, smaller language models, or ensembling methods. For example, Sun et al. (2019) uses a bidirectional Transformer to model the item sequence. Tang and Wang (2018) uses CNN to model user preference with convolutional filters.  Gu et al. (2020) and Yang et al. (2022) uses ensembling methods for recommendation.
B.2 Datasets and evaluating multimodal recommendation.
The development of multimodal recommendation systems has been facilitated by the availability of diverse datasets that incorporate various types of data, including text, images, and user interaction histories. Notable datasets in this domain include MovieLens-1M (Harper and Konstan, 19), which is extensively used for movie recommendations, and the Amazon dataset (Ni et al., 2019), which includes user reviews and metadata for product recommendations. For news and books recommendations, the MIND (Wu et al., 2020) and Goodreads (Wan and McAuley, 2018) datasets offer insights into user preferences in news and literature. Recently, LaMP (Salemi et al., 2023) introduces 7 text classification and text generation tasks across long contexts to evaluate LLM’s personalization capablity. In the context of outdoor activities, datasets such as Google Local Data (Li et al., 2022; Yan et al., 2023) and Yelp (Asghar, 2016) are crucial. These datasets not only provide textual reviews but also include user ratings and geospatial data, which are essential for recommending local businesses and services. These benchmarks utilize the traditional recommendation systems that focus on ranking user preferences based on historical IDs. Instead, we introduce new visual history benchmarks Google Review-V and Yelp-V, using data from Google Local and Yelp, to evaluate the multimodal recommendation systems under a more realistic visual history for task-agnostic setups.
C Prompt Template
We list the prompt template for aspect word generation in Figure 6, and the prompt template for candidate matching in Figure 7. The model will fill in category information in the [[Category]] slot and predict the answer at the [[Answer]] slot.
Figure 6: The prompt template for aspect word generation.
Figure 7: The prompt template for candidate matching.
D Additional Implementation Details
We present the hyperparameters for training VisualLens in PaliGemma and MiniCPM-V2.5 in Table 5 and Table 6. Note that since there are at most 100 candidates, we add special tokens <I1>, …, <I100> as identifier of the candidates in the vocab of the models. By doing so, we can predict the rank by taking the probability of only those special tokens in the first output token.
Hyperparameters for training on PaliGemma
Parameter Size 3B
Image Resolution 896 $\times$ 896
Number of Image Tokens 4096
Hidden Dimension Size 2048
LoRA Rank 16
LoRA $\alpha$ 16
LoRA dropout 0.1
GPU 8 $\times$ NVIDIA H100
Batch Size 8
Gradient Accumulation Steps 8
Warmup Steps 200
Learning Rate 1e-3
Table 5: Hyperparameters for training VisualLens with PaliGemma Backbone.
Hyperparameters for training on MiniCPM-V2.5
Parameter Size 8B
Image Resolution 980 $\times$ 980
Number of Image Tokens 96
Hidden Dimension Size 4096
LoRA Rank 64
LoRA $\alpha$ 64
LoRA dropout 0.1
GPU 8 $\times$ NVIDIA H100
Batch Size 8
Gradient Accumulation Steps 8
Warmup Steps 200
Learning Rate 1e-3
Table 6: Hyperparameters for training VisualLens with MiniCPM-V2.5 Backbone.